OpenStudy (anonymous):

When you solve the equation x + 9=15, how do you know what number to subtract from both sides? Why is it important to subtract the same number from both sides of the equation?

OpenStudy (jack1):

9 as you want the variable/unknown to exist on one side of the equals sign alone

OpenStudy (jack1):

pretend the equals sign is a seesaw, then visualise what you want to find, then manipulate that side of the equation so it is alone, and as its a seesaw, what you do to one side you must do to the other

OpenStudy (anonymous):

what?

OpenStudy (jack1):

x + 9 = 15 you want the x by itself, as it's the unknown factor so the equation you want to end up with is: x = something therefore the steps to get to that are: there is a +9 on the side with the x, so subtract 9 from both sides (otherwise it wont balance) left with: x +9 -9 = 15 -9 x = 6

OpenStudy (anonymous):

oh thx